Find a formula for the inverse of the function, if the function is one-to-one. If not, state that the function is not one-to-one for
F(x)=5x-3

Answer by CubeyThePenguin(3113)

Solve for x.

y = 5x - 3
y + 3 = 5x
(y + 3)/5 = x

f^(-1)(x) = (x + 3)/5